OpenStudy (anonymous):

Find the equation of both the tangent lines to the curve y=x^3-x that are parallel to the line 22x-2y+1=0

OpenStudy (anonymous):

is this geometry??

OpenStudy (anonymous):

No. It's calculus

OpenStudy (anonymous):

damm im not really familiar with it but i can try to hlp

OpenStudy (anonymous):

jus tell me ur main problem and ill see wat i can do

OpenStudy (anonymous):

Well, I'm sure to get the equation... I just don't know how to approach this question. I tried to differentiate both tangent lines...

OpenStudy (anonymous):

not sure *

OpenStudy (anonymous):

do u hve any notes that u can use??

OpenStudy (alekos):

how are you going with this? still need help?

OpenStudy (anonymous):

Yes, I still need help

OpenStudy (alekos):

so the 2 tangent lines must have the same slope as 22x-2y+1=0 do you know what the slope of this line is?

OpenStudy (anonymous):

So take the derivative of the 22x-2y+1=0?

OpenStudy (alekos):

No. first we re-arange the equation 22x+1=2y y=11x + 1/2 so what do you think the slope might be?

OpenStudy (anonymous):

11

OpenStudy (alekos):

that's it! so we're looking for two tangent lines with a slope of 11 i.e. y=11x+b now we take the derivative of y=x^3-x what do you think that would be?

OpenStudy (anonymous):

y'=3x^2-1

OpenStudy (alekos):

that's right so now we need to know at what values of x that this derivative (or slope of the tangent) is equal to 11 i.e. 3x^2 - 1 = 11

OpenStudy (alekos):

see if you can find the 2 values of x

OpenStudy (anonymous):

Is it +/-2?

OpenStudy (alekos):

very good. the tangent lines touch the cubic at x=2 & x=-2 so now we need to know the corresponding y-values using the original cubic equation. what would they be?

OpenStudy (anonymous):

Is it y=6 and y=-6?

OpenStudy (alekos):

well done! so know we can work out the two tangent equations y=11x+b by substituting these two points and evaluating b

OpenStudy (anonymous):

Oh I see! I think I got it from here!

OpenStudy (anonymous):

I got these two equations: 11x-y-16=0 and 11x-y+16=0

OpenStudy (alekos):

yep you got it. can also be written as y=11x+16 and y = 11x-16

OpenStudy (anonymous):

Okay! Thank you so much! :)

OpenStudy (alekos):

your welcome